(c) Show that for all values of θ the determinant







1 sin θ 1
− sin θ 1 sin θ
−1 − sin θ 1







lies between 2 and 4 inclusive. State one value of θ for which the determinant has the
value 2, and one for which it has the value 4.
1
Expert's answer
2020-07-07T20:09:24-0400

"\\begin{vmatrix} 1 & sin \\theta & 1\\\\ -sin \\theta & 1 & sin \\theta \\\\ -1& -sin \\theta & 1 \\end{vmatrix}\n\n\u200b\t\n \n\n\u200b"

"=1\cdot1\cdot1+(-1)\cdot sinθ \cdot sinθ +(- sinθ) \cdot (- sinθ) \cdot 1-"

"-(-1) \cdot 1 \cdot 1 - (-sinθ) \cdot sinθ \cdot 1-"

"- 1 \cdot (-sinθ) \cdot sinθ= 2+2(sinθ)^2"


"-1\le sinθ \le 1
\implies
 0\le (sinθ)^2 \le 1"

Then

"2 \le 2+2(sinθ)^2 \le 4"


If "\u03b8=0" then "2+2(sin\u03b8)^2\n =2"

If "\u03b8=\u03c0\/2" then "2+2(sin\u03b8)^2\n =4"